OpenStudy (anonymous):

|(8x+56)/56|>/10 I got -77 and 63 for the answers but I don't know how to put it in interval notation

OpenStudy (anonymous):

I tried every way and it says incorrect or i have overlapping intervals

OpenStudy (ashley1nonly):

Do you think that it is the right answer or do you need help solving

OpenStudy (ashley1nonly):

is the >10 or is it really >/10

OpenStudy (anonymous):

its > with a line under it

OpenStudy (anonymous):

Im 90 percent sure that is the right answer

OpenStudy (ashley1nonly):

let me check the work first okay

OpenStudy (anonymous):

okay

OpenStudy (ashley1nonly):

did you end up getting [-77,63]

OpenStudy (anonymous):

yea

OpenStudy (ashley1nonly):

did you have answer chocies

OpenStudy (anonymous):

what do you mean?

OpenStudy (ashley1nonly):

like is it multiple choice

OpenStudy (anonymous):

no

OpenStudy (ashley1nonly):

hmm sorry i am of no help but the only choice i would think it would be was [63,infinity). sorry

OpenStudy (anonymous):

its okay thanks

zepdrix (zepdrix):

Hey ms britt c:

OpenStudy (anonymous):

Hi again!

zepdrix (zepdrix):

\[\Large\rm \left|\frac{8x+56}{56}\right|\ge 10\]Multiply both sides by 56,\[\Large\rm \left|8x+56\right|\ge 560\]Then we end up with our two cases again when we drop the absolute bars.\[\Large\rm \cases{8x+56\ge560\\ \rm 8x+56\le-560}\]Solving for x in each case gives us:\[\Large\rm \cases{x\ge63\\ \rm x\le-77}\]So it looks like you've got the correct numbers. Just need some help writing in this in interval notation?

zepdrix (zepdrix):

Do you understand why the inequality flipped directions in the second case?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

dividing by a negative

zepdrix (zepdrix):

x greater than or equal to 63, So we'll have 63 and all the values larger than that heading towards positive infinity. x less than or equal to -77, We'll also have -77 and all the values smaller than that, heading towards negative infinity.

zepdrix (zepdrix):

Note: ALWAYS use a round bracket for \(\Large\rm \infty\). We don't "include" infinity in our interval since it doesn't represent an actual value. So for our lower side we'll have: \(\Large\rm (\infty,~ -77 \) What is different in this problem? What type of bracket do we want on the -77?

OpenStudy (anonymous):

]

zepdrix (zepdrix):

Good good, I just realized I made a typo, lemme fix the -infty,\[\Large\rm (-\infty,~ -77]\]So that's one of our pieces.

zepdrix (zepdrix):

How bout the other one, what do you think? c:

OpenStudy (anonymous):

(inf, 63]

OpenStudy (anonymous):

wait [63, inf)

zepdrix (zepdrix):

Yah there we go! We want the SMALLER number on the left since we read from left to right.\[\Large\rm [63,\infty)\]

zepdrix (zepdrix):

How do we put them together? Union or Intersection? Here, I'll give you a little cheat sheet.

zepdrix (zepdrix):